OpenStudy (fifciol):

I have to calculate the magnetic field of a square loop of wire at point P at height h above the center of loop. I get the book answer when I use small heights approximation (theta=alpha). Could someone tell me whether I'm right or wrong?

OpenStudy (fifciol):

 Book answer: \[B(h)=\frac{ 4\mu_0Ia^2 }{ \pi (4h^2+a^2)(4h^2+2a^2)^{\frac{ 1 }{ 2 }} }\]